Zeronknight (zeronknight):

Find the solution to the following inequality: x - 10 < 3 A. x < -7 B. x\leq -7 C. x\leq 13 D. x< 13

Zeronknight (zeronknight):

B \[x\leq -7\]

Zeronknight (zeronknight):

\[ C. x\leq 13\]

OpenStudy (anonymous):

What do you think the first step is in solving this inequality?

Zeronknight (zeronknight):

I know all the steps but I don't understand the underline under the < sign

OpenStudy (anonymous):

Ahh ok. That just means it is less than or equal to. For instance, if you had x<4, it would be all values less than 4 but not 4 itself. However, if you had: \[x \le 4\] It would be all numbers less than 4 as well as 4 itself.

Zeronknight (zeronknight):

I don't quite follow. So Every number is less than 4 ?

OpenStudy (anonymous):

Let's start with x<4. Do you understand what this is stating? This means that the value of x is everything less than 4, but nothing equal to 4 or greater than 4. Does that make sense?

Zeronknight (zeronknight):

Yes X < 4 basically means X is not equal and is less than 4. Stating every number before 4 is less than 4

Zeronknight (zeronknight):

but the underline < doesn't make any sense to me

OpenStudy (anonymous):

Right. So now, the underlined < sign just means you're also going to say that x=4 is true. So x<4 AND x=4. This means that everything less than 4 and 4 itself is the value of x. Does that help you understand?

Zeronknight (zeronknight):

So that means X also = 4. but X is also less than 4?

OpenStudy (anonymous):

Exactly. So x is all of the values less than 4 and x is the value 4 itself. They made it into a new symbol which combines the equals sign and the less than sign so they don't have to say x<4, x=4 every time.

Zeronknight (zeronknight):

-10 + 10 = 0. 7 + 10 = 17.

Zeronknight (zeronknight):

I forgot actually what am I supposed to do now?

OpenStudy (anonymous):

Which problem is that from?

Zeronknight (zeronknight):

The question I posted. Lol

OpenStudy (anonymous):

Ahh. Where did the 7 come from? I see: x-10 < 3

Zeronknight (zeronknight):

oh fail XD. I thought of 7 and wrote it...

Zeronknight (zeronknight):

I was looking at A,B, C and D

Zeronknight (zeronknight):

so basically 3 + 10 = 13 -10 + 10 = 0

OpenStudy (anonymous):

Right, so what inequality do you end up with?

Zeronknight (zeronknight):

See, i don't know if it's C or D

OpenStudy (anonymous):

Ahh. So you can never switch from a regular < or > to a \[\le or \ge\] during a problem like this. If you are finding the solution to the inequality, it will either keep the bar, or never have it throughout the problem.

Zeronknight (zeronknight):

ohh

Zeronknight (zeronknight):

So D

OpenStudy (anonymous):

Right.

Zeronknight (zeronknight):

Hello @mathmale

OpenStudy (mathmale):

A quick review: Find the solution to the following inequality: x - 10 < 3 Add 10 to both sides. Result: x<13. That's all. Finished. Supposing that you had \[x -3 \le 10\] to solve: Add 3 to both sides. Result: \[x \le13\] END!

OpenStudy (mathmale):

Hello, Zeronknight!

Zeronknight (zeronknight):

oh! Thank you for the explanation. I actually think I have a better understanding now. (I don't know if people would take this as sarcastic or not, but it's not sarcastic)

OpenStudy (mathmale):

My pleasure. Absolutely no suspicion of sarcasm! Just wanted you to know that the basic solution is quick and straightforward.

Zeronknight (zeronknight):

Just one question? When would I use the underlined <?

Zeronknight (zeronknight):

Well it would have it from the start right? But how would it exactly play in a problem?

OpenStudy (anonymous):

When it is equal to

Zeronknight (zeronknight):

Oh. So when X is equal to the number? I understands what it means, but not how it would right in a problem.

OpenStudy (mathmale):

< means "less than." \[\le .\] means "less than or equal to."

OpenStudy (mathmale):

You will encounter both in typical "inequality" problems. Your task is mainly to understand the difference between the two symbols and how to solve inequalities that involve them.

Zeronknight (zeronknight):

Oh

Zeronknight (zeronknight):

That makes much more sense, thank you.

OpenStudy (mathmale):

My great pleasure. Merry Christmas to you!

Zeronknight (zeronknight):

Merry Christmas too you too! Surprised anyone has said that. I usually encounter Happy Holidays XD

OpenStudy (mathmale):

;)

Zeronknight (zeronknight):

XD

OpenStudy (mathmale):

People try too hard to be politically correct. Every so often I'm politically incorrect. ;)

Zeronknight (zeronknight):

XD Well then. owo